. A consignment of 12 electronic components contains 1 component that is faulty. Two components are chosen randomly from this consignment for testing.

a. How many different combinations of 2 components could be chosen?

b. What is the probability that the faulty component will be chosen for testing?


1
Expert's answer
2022-04-29T14:44:27-0400

a.

"\dbinom{12}{2}=\dfrac{12!}{2!(12-2)!}=66"

b.


"\dfrac{\dbinom{1}{1}\dbinom{11}{1}}{\dbinom{12}{2}}=\dfrac{11}{66}=\dfrac{1}{6}"